# Nearest Neighbors GParareal: Improving Scalability of Gaussian Processes for Parallel-in-Time Solvers

### v1

@@ 5.3 Thomas labyrinth
where $a=0.5$ and $b=10$. We integrate it over $t\in[0,10]$ for $N=32$ and $N=64$, $t\in[0,40]$ for $N=128$, and $t\in[0,100]$ for $N=256$ and $N=512$ intervals. Following [22], we take $\boldsymbol{u}_{0}=(4.6722764,5.2437205e^{-10},-6.4444208e^{-10})$ as initial condition, such that the system exhibits chaotic dynamics. Further, we use RK1 with $10N$ steps for the coarse solver $\mathscr{G}$ and RK4 with $1e^{9}$ steps for the fine solver $\mathscr{F}$. We use the normalized version (see Step 1 of nnGParareal) with a normalized accuracy $\epsilon=5e^{-7}$. Also in this case, nnGParareal achieves the highest speed-up for all $N$ other than $N=64$, when $S_{\rm GPara}=3.79$ and $S_{{\rm nnGPara}}=3.65$ (we refer the reader to Table 8 in Supplementary Material D). Note also that the speed-ups for different $N$ are not comparable anymore, unless solving for the same $t_{N}$.
@@ 5.4 Viscous Burgers’ equation
Unlike for ODEs, nnGParareal requires more iterations than GParareal to converge when considering PDEs. This is also the case for the two-dimensional FitzHugh-Nagumo PDE in Section 5.5, suggesting a loss of information when considering the ($m$) nnGP instead of the GP trained on the entire dataset. Higher values of $m$ helps in reducing $K_{{\rm nnGPara}}$, lowering it down to 7 for $t_{N}=5$ and 12 for $t_{N}=5.9$ for $m$ as large as 30 (see Figure 15 in Supplementary Material F). At the same time though, the model cost increases too, so the highest speed-up may not be necessarily achieved by a larger $m$ and a lower $K_{{\rm nnGPara}}$ (see Figure 16 in Supplementary Material F). In general, even if $K_{{\rm nnGPara}}>K_{\rm GPara}$ for all considered $m$ values, the nnGParareal speed-up is always higher than that of the other algorithms, particularly when considering $t_{N}=5.9$, since (a) nnGParareal converges faster than Parareal for $t_{N}=5$ and much faster for $t_{N}=5.9$; (b) the high model cost of GParareal negatively offsets its gains in terms of $K_{\rm alg}$. As previously observed, Gparareal is an effective alternative to Pareal for moderate values of $N$, while the proposed nnGParareal makes it competitive and advantageous for large $N$ as well. Remarkably, both GParareal and nnGParareal are less sensible to the sub-optimal coarse solver chosen for $t_{N}=5.9$ than Parareal, with $K_{\rm GPara}$ and $K_{{\rm nnGPara}}$ rising slightly compared to $K_{\rm Para}$. This cushioning effect is particularly desired when in doubt about the accuracy of the selected $\mathscr{G}$.
@@ 5.5 FitzHugh-Nagumo PDE
As for the viscous Burgers’ PDE, if it terminates within the allocated 48 hours time budget, GParareal requires the smallest number of iterations to converge, followed by nnGParareal (see Figure 6, left panel). Nevertheless, except for $d=200$, when $\hat{S}_{\rm Para}>\hat{S}_{{\rm nnGPara}}$, nnGParareal outperforms the other two algorithms in terms of speed-up, as shown in Figure 6 (a detailed breakdown of all running times and training costs is reported in Table 9 in Supplementary Material E). The gap between theoretical and empirical speed-ups for nnGParareal can be closed using a more efficient parallel implementation (i.e. one where the overhead from parallelization is lower), while that, albeit small, between upper bound and theoretical speed-up by decreasing the model training cost (i.e. by using a cheaper model).
@@ 6 Discussion
The idea of nnGParareal is to replace the GP trained on the entire dataset with one trained on $m$ data points. Among several approaches, we numerically illustrate how choosing $m$ as the nearest neighbors leads to the highest prediction accuracy and the smaller number of iterations needed to converge. By embedding the derived nnGP within Parareal, we obtain an algorithm which mostly preserves or even improves the accuracy of GParareal (measured in terms of number of iterations needed to converge), thanks to the partial relaxation of the second-order stationarity assumption made by the GPs [10], while drastically reducing its computational cost and wallclock time, thus improving the speed-up with respect to running $\mathscr{F}$ sequentially. Indeed, the time complexity of nnGParareal is now only $O(dm^{3}N+dN\log(kN))$ at iteration $k$ instead of $O(d(kN)^{3})$ of GParareal. This allows nnGParareal to tackle GParareal’s scalability issues, consistently yielding the highest speed-up compared to GParareal and Parareal across several different systems of ODEs and PDEs. This makes our algorithm better suited to the modern requirements and capabilities of HPC, thus expanding the temporal scale over which we can solve ODEs and PDEs with time parallelization. Nevertheless, the performance of nnGParareal might be further improved by leaving the Gaussian processes framework and selecting a learned model capable of training on all $d$ coordinates simultaneously, with a cheaper hyperparameter optimization routine. We leave this for future work.
@@ A.4 Double pendulum
with $f_{1}\left(u_{1},u_{2}\right)=\sin\left(u_{1}-u_{2}\right)\cos\left(u_{1}-u_{2}\right)$ and $f_{2}\left(u_{1},u_{2}\right)=2-\cos^{2}\left(u_{1}-u_{2}\right)$ and where $m,\ell$, and $g$ have been scaled out of the system by letting $\ell=g$. The variables $u_{1}$ and $u_{2}$ measure the angles between each pendulum and the vertical axis, while $u_{3}$ and $u_{4}$ measure the corresponding angular velocities. The system exhibits chaotic behavior and is commonly used in the literature. Based on the initial condition, it can be difficult to learn. We integrate over $t\in[0,80]$ using $N=32$ intervals, taking $\boldsymbol{u}_{0}=(-0.5,0,0,0)$ as initial condition. We use RK1 with $3104$ steps for the coarse solver $\mathscr{G}$, and RK8 with $2.17e^{5}$ steps for the fine solver $\mathscr{F}$. This is a similar setting as [44, Figure 4.10], although, like above, we use the normalized version and set a normalized $\epsilon=5e^{-7}$.
@@ A.6 FitzHugh-Nagumo PDE
(25) $\mathcal{K}_{\rm GP}(\widetilde{\boldsymbol{U}}_{i}^{k},\widetilde{\boldsymbol{U}}_{i^{\prime}}^{k^{\prime}})$ $=\mathcal{K}_{\rm GP}\left((\boldsymbol{U}_{i}^{k},i,k),(\boldsymbol{U}_{i^{\prime}}^{k^{\prime}},i^{\prime},k^{\prime})\right)$
@@
where $\sigma_{\rm o}^{2}$ is the output length scale, as in (6). Collectively, $\boldsymbol{\theta}=(\sigma_{\rm o}^{2},\sigma_{s}^{2},\sigma_{i}^{2},\sigma_{k}^{2},\sigma_{\rm reg}^{2})$ form the hyper-parameters of the GP augmented to include time information. When combined with Parareal and nn training, we refer to this as nnGParareal with Time.
@@
We aim to change the choice of the $m$ neighbors to account for time. Since the kernel estimates the correlation between observations, it seems natural to build $\widetilde{D}_{i-1,k}$ for a prediction $\widehat{f}(\widetilde{\boldsymbol{U}}_{i-1}^{k})$ out of the $m$ pairs $(\widetilde{\boldsymbol{U}},(\mathscr{F}-\mathscr{G})(\widetilde{\boldsymbol{U}}))$ which have the highest kernel score $\mathcal{K}_{\rm GP}(\widetilde{\boldsymbol{U}},\widetilde{\boldsymbol{U}}_{i-1}^{k})$, $\widetilde{\boldsymbol{U}}\in\widetilde{\mathcal{U}}_{k}$. This was already identified by [59] as a natural strategy for subset selection and is also what is used in[12]. However, as pointed out in the same paper, the ranking operation requires knowledge of the optimal hyper-parameters $\boldsymbol{\theta}^{*}$ to evaluate the kernel $\mathcal{K}_{\rm GP}(\widetilde{\boldsymbol{U}},\widetilde{\boldsymbol{U}}_{i-1}^{k})$. In general, $\boldsymbol{\theta}^{*}$ is not available unless a data subset $\widetilde{\mathcal{D}}_{i,k}$ is proposed first, and the marginal log-likelihood is minimized for $\boldsymbol{\theta}$. This suggests a computationally expensive iterative procedure. Propose $\widetilde{\mathcal{D}}_{i-1,k}^{0}$ at random and estimate $\boldsymbol{\theta}_{0}^{*}$ conditional on $\widetilde{\mathcal{D}}_{i-1,k}^{0}$. Using $\boldsymbol{\theta}^{*}_{0}$, rank the observation in $\widetilde{D}_{k}$ to obtain a new subset $\widetilde{\mathcal{D}}_{i-1,k}^{1}$. Estimate the new optimal hyper-parameters $\boldsymbol{\theta}_{1}^{*}$ based on $\widetilde{\mathcal{D}}_{i-1,k}^{1}$. Repeat until the data subset has stabilized or for an arbitrary number of iterations. Since each iteration requires optimizing $\boldsymbol{\theta}$, which is usually expensive, such a strategy has historically been discarded in favor of more straightforward rules. For example, [10] show empirically that simply choosing the $m$ nearest neighbors gives excellent performance compared to the full GP. Furthermore, [11] states: “As nearest neighbors correspond to points with highest spatial correlation, this choice produces excellent approximations.” We observed in practice that two or three iterations of the procedure are sufficient for the subsets $\widetilde{D}_{i-1,k}^{\cdot}$ to stabilize in most cases. Moreover, the converged subset consisted of the $m$ closest spatial neighbors in over 80% of the cases, as seen Figure 14 for the Lorenz system. Finally, to verify that the nnGParareal with Time does not provide further performance gains over nnGParareal, we compare the two over a range of systems in Table 7. Since they show equivalent performance, we prefer the simpler nn variant, which is much cheaper to run.
@@
Supplementary Material D Thomas Labyrinth Scalability Table
@@
Supplementary Material E Scalability results for the FHN PDE model
@@
Supplementary Material F Burgers’ performance across $m$
This section reports additional experimental results for the Burgers’ equation in Section 5.4. In particular, we study the impact of different dataset sizes $m$ on nnGParareal performance. Analogous analysis has been carried out for other systems, such as Thomas labyrinth and FitzHugh-Nagumo PDE, which yielded similar results and hence are not reported here. For both $T=5$ and $T=5.9$, Figure 15 reports the histogram of the iterations to convergence across several choices of $m$ ranging from 11 to 30. For each $m$, we executed 100 independent runs of nnGParareal to account for the algorithm’s randomness coming from the log-likelihood exploration (see Supplementary Material G). The performance is largely unaffected by changes in $m$, with minor gains resulting from higher $m$, as could be expected. Figure 16 displays the empirical distribution of the speed-up corresponding to the simulations in Figure 15. Although nnGParareal takes longer to converge than GParareal for Burgers’ equation, on average, the speed-up gains are always greater due to the lower training cost of the model.
@@
Supplementary Material G Jitter and hyper-parameters optimization
@@
We found the tuning of $\sigma_{\rm reg}^{2}$ to be important for (nn)GParareal, given the close distance between points in the dataset, as seen in Figure 13. This is a situation prone to numerical instability [38]. Some authors recommend setting a lower bound for the jitter to ensure the matrix does not leave the symmetric positive definite cone. These values are either found analytically [43] or via experiments [48]. Others [7, 38] advocate for the use of maximum likelihood (MLE) or leave-one-out cross-validation (LOO-CV), which can be computed in closed-form, for a data-driven tuning approach. In our application, we found MLE slightly more stable than LOO-CV in selecting $\sigma_{\rm reg}^{2}$, which is searched on a grid of size $n_{\rm reg}$, with values as low as -20 on a log scale. All GParareal and nnGParareal results in this work feature automatic nugget selection. The remaining hyper-parameters $(\sigma^{2}_{\rm i},\sigma^{2}_{\rm o})$ are also tuned by MLE, using the Nelder-Mead optimization algorithm. Evidence of the advantages of MLE in the scope of hyper-parameters optimization is shown in [47]. Local minima are addressed using random restarts, i.e. by running the optimization procedure Nelder-Mead $n_{\rm start}$ times, independently, with different random initial guesses for the hyperparameters $\boldsymbol{\theta}$.

### v2

@@ 5.3 Thomas labyrinth
where $a=0.5$ and $b=10$. We integrate it over $t\in[0,10]$ for $N=32$ and $N=64$, $t\in[0,40]$ for $N=128$, and $t\in[0,100]$ for $N=256$ and $N=512$ intervals. Following [23], we take $\boldsymbol{u}_{0}=(4.6722764,5.2437205e^{-10},-6.4444208e^{-10})$ as initial condition, such that the system exhibits chaotic dynamics. Further, we use RK1 with $10N$ steps for the coarse solver $\mathscr{G}$ and RK4 with $1e^{9}$ steps for the fine solver $\mathscr{F}$. We use the normalized version (see Step 1 of nnGParareal) with a normalized accuracy $\epsilon=5e^{-7}$. Also in this case, nnGParareal achieves the highest speed-up for all $N$ other than $N=64$, when $S_{\rm GPara}=3.79$ and $S_{{\rm nnGPara}}=3.65$ (we refer the reader to Table 8 in Supplementary Material E). Note also that the speed-ups for different $N$ are not comparable anymore, unless solving for the same $t_{N}$.
@@ 5.4 Viscous Burgers’ equation
Unlike for ODEs, nnGParareal requires more iterations than GParareal to converge when considering PDEs. This is also the case for the two-dimensional FitzHugh-Nagumo PDE in Section 5.5, suggesting a loss of information when considering the ($m$) nnGP instead of the GP trained on the entire dataset. Higher values of $m$ helps in reducing $K_{{\rm nnGPara}}$, lowering it down to 7 for $t_{N}=5$ and 12 for $t_{N}=5.9$ for $m$ as large as 30 (see Figure 16 in Supplementary Material G). At the same time though, the model cost increases too, so the highest speed-up may not be necessarily achieved by a larger $m$ and a lower $K_{{\rm nnGPara}}$ (see Figure 17 in Supplementary Material G). In general, even if $K_{{\rm nnGPara}}>K_{\rm GPara}$ for all considered $m$ values, the nnGParareal speed-up is always higher than that of the other algorithms, particularly when considering $t_{N}=5.9$, since (a) nnGParareal converges faster than Parareal for $t_{N}=5$ and much faster for $t_{N}=5.9$; (b) the high model cost of GParareal negatively offsets its gains in terms of $K_{\rm alg}$. As previously observed, Gparareal is an effective alternative to Parareal for moderate values of $N$, while the proposed nnGParareal makes it competitive and advantageous for large $N$ as well. Remarkably, both GParareal and nnGParareal are less sensible to the sub-optimal coarse solver chosen for $t_{N}=5.9$ than Parareal, with $K_{\rm GPara}$ and $K_{{\rm nnGPara}}$ rising slightly compared to $K_{\rm Para}$. This cushioning effect is particularly desired when in doubt about the accuracy of the selected $\mathscr{G}$.
We conclude this section by investigating whether the performance observed for Parareal and nnGParareal may result from asymmetries in the finer temporal/spatial resolutions, as an unbalanced discretization may “artificially”enhance speed-up gains. To assess whether this is the case, we perform two additional simulations. First, we increase the spatial discretization resolution to $d=400,815,1128$, while simultaneously reducing the temporal resolution of the fine solver $\mathcal{F}$ to $10^{4},10^{5},$ and $6\times 10^{5}$ steps, respectively. In all tested scenarios, nnGParareal consistently achieves greater speed-ups than Parareal (results not shown), mirroring the findings summarized in Table 4. Although this adjustment partially reduces the differences in orders of magnitude of the temporal and spatial discretization errors, it still does not quantify whether the two errors are effectively balanced. Direct quantification and balancing of spatial and temporal discretization errors (error equilibration) are inherently challenging tasks in complex high-dimensional PDE problems, like those considered here. Without access to analytical solutions, error estimation typically demands highly accurate numerical approximations in both time and space, making it computationally infeasible. To overcome this limitation, we conduct an additional numerical experiment using the heat equation, a PDE with an explicitly known analytical solution, enabling the derivation of balanced discretization errors. For this experiment, we choose the initial condition $u_{0}(x)=\sin(2\pi x)$, and evaluate the performance of both Parareal and nnGParareal in a setting specifically designed to ensure balanced discretization errors, benchmarking their results against a sequential exact solver as baseline. This experiment, detailed in Supplementary Material H, demonstrates parallel speed-ups of $8.85$ for Parareal and $23.50$ for nnGParareal.
@@ 5.5 FitzHugh-Nagumo PDE
As for the viscous Burgers’ PDE, if it terminates within the allocated 48 hours time budget, GParareal requires the smallest number of iterations to converge, followed by nnGParareal (see Figure 6 , left panel). Nevertheless, except for $d=200$, when $\hat{S}_{\rm Para}>\hat{S}_{{\rm nnGPara}}$, nnGParareal outperforms the other two algorithms in terms of speed-up, as shown in Figure 6 (a detailed breakdown of all running times and training costs is reported in Table 9 in Supplementary Material F). The gap between theoretical and empirical speed-ups for nnGParareal can be closed using a more efficient parallel implementation (i.e. one where the overhead from parallelization is lower), while that, albeit small, between upper bound and theoretical speed-up by decreasing the model training cost (i.e. by using a cheaper model).
@@ 6 Discussion
The idea of nnGParareal is to replace the GP trained on the entire dataset with one trained on $m$ data points. Among several approaches, we numerically illustrate how choosing $m$ as the nearest neighbors leads to the highest prediction accuracy and the smaller number of iterations needed to converge. By embedding the derived nnGP within Parareal, we obtain an algorithm which mostly preserves or even improves the accuracy of GParareal (measured in terms of number of iterations needed to converge), thanks to the partial relaxation of the second-order stationarity assumption made by the GPs [10], while drastically reducing its computational cost and wallclock time, thus improving the speed-up with respect to running $\mathscr{F}$ sequentially. Indeed, the time complexity of nnGParareal is now only $O(dm^{3}N+dN\log(kN))$ at iteration $k$ instead of $O(d(kN)^{3})$ of GParareal. This allows nnGParareal to tackle GParareal’s scalability issues, consistently yielding the highest speed-up compared to GParareal and Parareal across several different systems of ODEs and PDEs. This makes our algorithm better suited to the modern requirements and capabilities of HPC, thus expanding the temporal scale over which we can solve ODEs and PDEs with time parallelization. Nevertheless, the performance of nnGParareal might be further improved by leaving the Gaussian processes framework and selecting a learned model capable of training on all $d$ coordinates simultaneously, with a cheaper hyperparameter optimization routine, as recently proposed in [22] via random weights neural networks.
Finally, it is worth recalling that both GParareal and nnGParareal use only the underlying GP and nnGP posterior means for prediction, ignoring the underlying uncertainty in the GP and nnGP posteriors. Propagating the full uncertainty will yield a probabilistic version of GParareal and nnGParareal, respectively, as discussed in [45]. We leave this for future work.
@@ A.4 Double pendulum
with $f_{1}\left(u_{1},u_{2}\right)=\sin\left(u_{1}-u_{2}\right)\cos\left(u_{1}-u_{2}\right)$ and $f_{2}\left(u_{1},u_{2}\right)=2-\cos^{2}\left(u_{1}-u_{2}\right)$ and where $m,\ell$, and $g$ have been scaled out of the system by letting $\ell=g$. The variables $u_{1}$ and $u_{2}$ measure the angles between each pendulum and the vertical axis, while $u_{3}$ and $u_{4}$ measure the corresponding angular velocities. The system exhibits chaotic behavior and is commonly used in the literature. Based on the initial condition, it can be difficult to learn. We integrate over $t\in[0,80]$ using $N=32$ intervals, taking $\boldsymbol{u}_{0}=(-0.5,0,0,0)$ as initial condition. We use RK1 with $3104$ steps for the coarse solver $\mathscr{G}$, and RK8 with $2.17e^{5}$ steps for the fine solver $\mathscr{F}$. This is a similar setting as [45, Figure 4.10], although, like above, we use the normalized version and set a normalized $\epsilon=5e^{-7}$.
@@ A.6 FitzHugh-Nagumo PDE
$\mathcal{K}_{\rm GP}(\widetilde{\boldsymbol{U}}_{i}^{k},\widetilde{\boldsymbol{U}}_{i^{\prime}}^{k^{\prime}})$ $=\mathcal{K}_{\rm GP}\left((\boldsymbol{U}_{i}^{k},i,k),(\boldsymbol{U}_{i^{\prime}}^{k^{\prime}},i^{\prime},k^{\prime})\right)$
@@
where $\sigma_{\rm o}^{2}$ is the output length scale, as in (6). Collectively, $\boldsymbol{\theta}=(\sigma_{\rm o}^{2},\sigma_{s}^{2},\sigma_{i}^{2},\sigma_{k}^{2},\sigma_{\rm reg}^{2})$ form the hyperparameters of the GP augmented to include time information. When combined with Parareal and nn training, we refer to this as nnGParareal with Time.
@@
We aim to change the choice of the $m$ neighbors to account for time. Since the kernel estimates the correlation between observations, it seems natural to build $\widetilde{D}_{i-1,k}$ for a prediction $\widehat{f}(\widetilde{\boldsymbol{U}}_{i-1}^{k})$ out of the $m$ pairs $(\widetilde{\boldsymbol{U}},(\mathscr{F}-\mathscr{G})(\widetilde{\boldsymbol{U}}))$ which have the highest kernel score $\mathcal{K}_{\rm GP}(\widetilde{\boldsymbol{U}},\widetilde{\boldsymbol{U}}_{i-1}^{k})$, $\widetilde{\boldsymbol{U}}\in\widetilde{\mathcal{U}}_{k}$. This was already identified by [60] as a natural strategy for subset selection and is also what is used in[12]. However, as pointed out in the same paper, the ranking operation requires knowledge of the optimal hyperparameters $\boldsymbol{\theta}^{*}$ to evaluate the kernel $\mathcal{K}_{\rm GP}(\widetilde{\boldsymbol{U}},\widetilde{\boldsymbol{U}}_{i-1}^{k})$. In general, $\boldsymbol{\theta}^{*}$ is not available unless a data subset $\widetilde{\mathcal{D}}_{i,k}$ is proposed first, and the marginal log-likelihood is minimized for $\boldsymbol{\theta}$. This suggests a computationally expensive iterative procedure. Propose $\widetilde{\mathcal{D}}_{i-1,k}^{0}$ at random and estimate $\boldsymbol{\theta}_{0}^{*}$ conditional on $\widetilde{\mathcal{D}}_{i-1,k}^{0}$. Using $\boldsymbol{\theta}^{*}_{0}$, rank the observation in $\widetilde{D}_{k}$ to obtain a new subset $\widetilde{\mathcal{D}}_{i-1,k}^{1}$. Estimate the new optimal hyperparameters $\boldsymbol{\theta}_{1}^{*}$ based on $\widetilde{\mathcal{D}}_{i-1,k}^{1}$. Repeat until the data subset has stabilized or for an arbitrary number of iterations. Since each iteration requires optimizing $\boldsymbol{\theta}$, which is usually expensive, such a strategy has historically been discarded in favor of more straightforward rules. For example, [10] show empirically that simply choosing the $m$ nearest neighbors gives excellent performance compared to the full GP. Furthermore, [11] states: “As nearest neighbors correspond to points with highest spatial correlation, this choice produces excellent approximations.” We observed in practice that two or three iterations of the procedure are sufficient for the subsets $\widetilde{D}_{i-1,k}^{\cdot}$ to stabilize in most cases. Moreover, the converged subset consisted of the $m$ closest spatial neighbors in over 80% of the cases, as seen Figure 14 for the Lorenz system. Finally, to verify that the nnGParareal with Time does not provide further performance gains over nnGParareal, we compare the two over a range of systems in Table 7. Since they show equivalent performance, we prefer the simpler nn variant, which is much cheaper to run.
@@
Figure 15: Effect of increased coarse solver precision (i.e., number of steps) on both Parareal and nnGParareal speed-up, for the Hopf bifurcation system with $N=32$ and $1.74e^{8}$ fine solver steps. Parareal fails to converge for $10^{3}$ steps (missing data in the plot), while nnGParareal converges with a notable speed-up.
Supplementary Material D Hopf bifurcation’s performance across $N_{\mathcal{G}}$
Figure 15 shows the speed-ups of Parareal and nnGParareal as a function of the number of corse solver (RK1) steps for the Hopf bifurcation system with $N=32$ and $1.74e^{8}$ steps of the fine solver (RK8). While the speed-ups eventually become similar for more than $10^{5}$ coarse steps, overall nnGParareal allows to use much cheaper coarse solver precision than Parareal. Indeed, the latter requires many more coarse steps than nnGParareal to obtain a similar speed-up (e.g., $10^{4}$ vs $10^{3}$ for a speed-up of 50), and does not converge for $10^{3}$ coarse solver steps.
Supplementary Material E Thomas Labyrinth Scalability Table
@@
Supplementary Material F Scalability results for the FHN PDE model
@@
Supplementary Material G Burgers’ performance across $m$
This section reports additional experimental results for the Burgers’ equation in Section 5.4. In particular, we study the impact of different dataset sizes $m$ on nnGParareal performance. Analogous analysis has been carried out for other systems, such as Thomas labyrinth and FitzHugh-Nagumo PDE, which yielded similar results and hence are not reported here. For both $T=5$ and $T=5.9$, Figure 16 reports the histogram of the iterations to convergence across several choices of $m$ ranging from 11 to 30. For each $m$, we executed 100 independent runs of nnGParareal to account for the algorithm’s randomness coming from the log-likelihood exploration (see Supplementary Material I). The performance is largely unaffected by changes in $m$, with minor gains resulting from higher $m$, as could be expected. Figure 17 displays the empirical distribution of the speed-up corresponding to the simulations in Figure 16. Although nnGParareal takes longer to converge than GParareal for Burgers’ equation, on average, the speed-up gains are always greater due to the lower training cost of the model.
@@
Supplementary Material H Heat equation
We consider the one‐dimensional heat equation on the spatial interval $x\in(0,L)$ over time $t\in(0,T]$,
$\frac{\partial u}{\partial t}(x,t)\;=\;\alpha\,\frac{\partial^{2}u}{\partial x^{2}}(x,t),$
subject to homogeneous Dirichlet boundary conditions
$u(0,t)=u(L,t)=0,\quad t\in(0,T],$
and initial condition $u(x,0)=u_{0}(x)$. The heat equation describes, among other phenomena, heat distribution in a homogeneous isotropic medium, where $u(x,t)$ represents temperature at position $x$ and time $t$, and $\alpha$ denotes thermal diffusivity. We discretize the one dimensional spatial domain using finite difference [17], defining $d+1$ equally spaced points $x_{j+1}=x_{j}+\Delta x$, with $\Delta x=L/d$, $j=0,\ldots,d$, resulting in a $d$-dimensional ODE system. We choose parameters $L=1$, $T=2$, $\alpha=0.1$, $d=40$, and $u_{0}(x)=\sin(2\pi x)$ for our numerical experiments. Given the periodic initial condition, the exact solution is
(23) $u(x,t)=\exp\bigl{(}-0.4\pi^{2}\,t\bigr{)}\,\sin(2\pi x).$
Parameter Description
$N$ Number of Parareal intervals 300
-- Number of processors for parallelization (set equal to $N$) 300
$\mathscr{F}$ Fine solver Runge-Kutta 8
$\mathscr{G}$ Coarse solver Runge-Kutta 2
$N_{\mathscr{F}}$ Number of temporal discretization steps per interval for $\mathscr{F}$ 1000
$N_{\mathscr{G}}$ Number of temporal discretization steps per interval for $\mathscr{G}$ 2
$N\times N_{\mathscr{G}}$ Total number of temporal steps for $\mathscr{G}$ across all intervals 600
Table 10: Parareal simulation parameters
Table 10 summarizes the Parareal simulation parameters. To ensure balanced discretization errors, we estimate temporal and spatial errors of the fine sequential solver $\mathscr{F}$ independently, proceeding as detailed before:
  • Increase spatial (resp. temporal) resolution by a factor of $15$.
  • Keeping temporal (resp. spatial) resolution fixed, numerically approximate the solution.
  • Measure temporal (resp. spatial) discretization errors, $e_{t}$ (resp. $e_{s}$), using the Euclidean distance to the exact solution (23).
With this approach, we obtained time and space discretization errors equal to $e_{s}=2.13\times 10^{-4}$ and $e_{t}=8.32\times 10^{-4}$, respectively, confirming that both errors are indeed of the same order of magnitude and balanced. Using this balanced resolution, we compared the runtimes of Parareal and nnGParareal against those of the sequential fine solver. The algorithms demonstrated significant speedups even with a minimal number of coarse solver steps, as shown in Table 11.
Algorithm $K_{\rm alg}$ $T_{\mathscr{G}}$ $T_{\mathscr{F}}$ $T_{\rm model}$ $T_{\rm alg}$ $\widehat{S}_{\rm alg}$
Fine -- -- -- -- 98.93 1
Parareal 33 0.16 0.33 0.003 11.18 8.85
nnGParareal 3 0.13 0.37 2.95 4.20 23.50
Table 11: Performance of Parareal and nnGParareal ($m=18$) for the heat equation (23). Times are in seconds. $T_{\mathscr{F}}$ and $T_{\mathscr{G}}$ represent the fine/coarse solvers’ runtime per iteration, $K_{\rm alg}$ is iteration count to convergence, $T_{\rm model}$ includes training/hyperparameter selection/prediction time, and $T_{\rm alg}$ and $\widehat{S}_{\rm alg}$ represent total runtime and empirical parallel speedup, respectively.
While we acknowledge the moderate spatial dimension and relatively low time resolution, which typically do not warrant parallelization across $N=300$ processors, this controlled example demonstrates that our method, and PinT methods more broadly, do not rely on significant over-resolution in one discretization dimension to achieve meaningful parallel performance.
Supplementary Material I Jitter and hyperparameters optimization
@@
We found the tuning of $\sigma_{\rm reg}^{2}$ to be important for (nn)GParareal, given the close distance between points in the dataset, as seen in Figure 13. This is a situation prone to numerical instability [39]. Some authors recommend setting a lower bound for the jitter to ensure the matrix does not leave the symmetric positive definite cone. These values are either found analytically [44] or via experiments [49]. Others [7, 39] advocate for the use of maximum likelihood (MLE) or leave-one-out cross-validation (LOO-CV), which can be computed in closed-form, for a data-driven tuning approach. In our application, we found MLE slightly more stable than LOO-CV in selecting $\sigma_{\rm reg}^{2}$, which is searched on a grid of size $n_{\rm reg}$, with values as low as -20 on a log scale. All GParareal and nnGParareal results in this work feature automatic nugget selection. The remaining hyperparameters $(\sigma^{2}_{\rm i},\sigma^{2}_{\rm o})$ are also tuned by MLE, using the Nelder-Mead optimization algorithm. Evidence of the advantages of MLE in the scope of hyperparameters optimization is shown in [48]. Local minima are addressed using random restarts, i.e. by running the optimization procedure Nelder-Mead $n_{\rm start}$ times, independently, with different random initial guesses for the hyperparameters $\boldsymbol{\theta}$. Optimal hyperparameters computed during previous iterations can also be used as starting points for the optimization procedure.